A person throws a penny off the roof of a building with an initial velocity of -10m/s (she throws it down). If the roof is 125 m above the ground, determine how fast the penny will be going when it reaches street level
1
Expert's answer
2020-09-22T15:30:42-0400

We can find the velocity of the penny when it reaches street level from the kinematic equation (also, we choose the upwards as the positive direction):


"v_f^2 = v_i^2 + 2ad,"

here, "v_f" is the final velocity of the penny when it reaches street level, "v_i = -10 \ ms^{-1}" is the initial velocity of the penny, "a = g = -9.8 \ ms^{-2}" is the acceleration due to gravity (since we choose the upwards as the positive direction it will be with the sign minus), "d = -125 \ m" is the displacement of the penny (or the height of the roof above the ground).

Then, we get:


"v_f = \sqrt{v_i^2 + 2ad},""v_f = \sqrt{(-10 \ ms^{-1})^2 + 2 \cdot (-9.8 \ ms^{-2}) \cdot (-125 \ m)} = 50.5 \ ms^{-1}."

Answer:

"v_f = 50.5 \ ms^{-1}."